1 -2/3y=9/5 -y/5 +3/5
Answer by stanbon(75887)

1 -(2/3)y = (9/5) -(y/5) +(3/5)
-------------------
Multiply thru by 15 to get rid of the fractions:
15 - 30y = 27 - 3y + 9
Simplify:
15 - 30y = 36 - 3y
27y = -21
y = -7/9
=================
Cheers,
Stan H.